OpenStudy (natsent06):

Identify the common factor in -8e + 9e.

OpenStudy (ahsome):

The common factor is \(e\) Divide both by \(e\) \[\frac{-8e+9e}{e}=-8+9\] No other factors from there

OpenStudy (natsent06):

so it would be

OpenStudy (ahsome):

If this helped you, please press the "Best Response" and "Close Question" so others can see your question is solved. Thanks :)

OpenStudy (natsent06):

@Ahsome what is the answer

OpenStudy (ahsome):

The common factor is \(e\)